Please help me in this question as soon as possible. thank you
 
1/x+1/y+1/z = 1 where x is less than y and y is less than z. Find x,y,z
 
 
 

Grade:

2 Answers

Sahil Shah
8 years ago
x=2, y=3, z=6
I do not know how to get to this. I just substituted these values.
I think these conditions must be fulfilled:
z=x*y
x+y=z-1
 
Sahil Shah
8 years ago
x=2, y=3, z=6
I do not know how to get to this. I just substituted these values.
I think these conditions must be fulfilled:
z=x*y
x+y=z-1